I've tried the problems like these a million times and always come out with a really off answer.
(-2x) + (-x)
some of the problems are more complex:
(-4) + 2x + 2x

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(48575)

(-2x) + (-x) = -3x
----------------------------------
some of the problems are more complex:
(-4) + 2x + 2x = 0
----------------------------------
Cheers,
Stan H.

Answer by Alan3354(21612)

I've tried the problems like these a million times and always come out with a really off answer.
(-2x) + (-x) = -3x
---------
some of the problems are more complex:
(-4) + 2x + 2x = 4x - 4